9-5n+1=15. What does n equal?
Answer by Cromlix(4381)

9-5n+1=15
Collect n values on left and
numbers on right.
Don't forget to change the value's
sign if you are changing its side.
-5n = 15 - 9 -1
-5n = 5
n = -1
Hope this helps
:-)